40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Mar 10 11:46:22 2021, Paco, Anchal, Summary, IMC, IMC free swinging prep 
    Reply  Wed Mar 10 15:00:16 2021, gautam, Summary, IMC, IMC free swinging prep 
       Reply  Wed Mar 10 15:29:58 2021, Anchal, Summary, IMC, IMC free swinging prep 
          Reply  Thu Jun 3 17:35:31 2021, Anchal, Summary, IMC, Fixed medm button C1IOO_Mech_Shutters.png
    Reply  Wed Mar 10 15:35:25 2021, Paco, Anchal, Summary, IMC, IMC free swinging experiment set to trigger at 5:00 am freeSwingMC.py.zip
Message ID: 15893     Entry time: Wed Mar 10 11:46:22 2021     Reply to this: 15895   15897
Author: Paco, Anchal 
Type: Summary 
Category: IMC 
Subject: IMC free swinging prep 

[Paco, Anchal]

# Initial State
- MC is locked. The PRM monitor shows some oscillations.
- POP monitor shows light flashing once in a while.
- AS monitor shows one beam along with some other flashing beam around it.
- PRM Watchdog is tripped and shutdown. Everything else is normal except for overload on SRM OpLevs.
- Donatella got a mouse promotion

# Reenabling PRM watchdog:
- The custom reEnablePRMWatchdog.py has been deleted.
- Tried enabling the coil outputs manually and switching watchdog to Normal.
- Again saw large fluctuations like yesterday.
- Probably still the same issue of how current calculated actuations to the coils is in range -600 to -900 and gives and impulse to the optics when suddenly turned on.
- Waiting for PRM to damp down a little.
- Today we plan to change the position bias on PRM C1:SUS-PRM_POS_OFFSET instead of changing biases in pitch and yaw.
- Changing C1:SUS-PRM_POS_OFFSET from 0 to +/- 100 without enabling the coils, it seems upper and lower coils are anticorrelated with just changing the position. So going back to changing pitch.
- Changing C1:SUS-PRM_PIT_OFFSET from 0 -> 780. Switched on watchdog to normal.
- PRM damped down. OpLev errors are also within range.
- Enabled both OpLevs.

# Try locking Y-Arm
- IFO>CONFIGURE>YARM>Restore YARM (POY) using Donatella. See a bunch of python error messages in the call complaining about unable to find some python 2 files. Closed it with Ctrl-C after a stuck state.
- Tried running it on Pianosa, the script ran without error but Y-Arm didn't lock.

# Try locking X-Arm
- IFO>CONFIGURE>XARM>Restore XARM (POX) on Donatella. Again a bunch of OSError messages. Donatella is not configured properly to run scripts.
- Tried running it on Piasnosa, the script ran without error but X-Arm didn't lock.
- This might mean that both arms are misaligned or the BS/PRM is misaligned.
- Moving around C1:SUS-PRM_PIT_OFFSET and C1:SUS-PRM_YAW_OFFSET in order to see if the transmitted light is misalgined. Both arms are set to acquire lock if possible. No luck.

# Hypothesis: The Arm cavity is not aligned within itself (ITM-ETM)
- Will try to lock X-Arm with green light while tuning the ETMX. Hopefully the BS and ITM are aligned so that once we align ETMX to get a green lock, the IR will also lock from the other side.
- Running IFO>CONFIGURE>XARM>Restore XARM (ALS) on Pianosa. No lock, moving forward with tunning ETMX pitch and yaw offsets. Nothing changed. Brought back to same values.

[Rana joined, Anchal moved to Rossa from Pianosa]

# Moving on to IMC suspensions characterization:
- Closed the PSL shutter, to our suprise, the MC was still locked. We thought this would take away any light from IMC but it doesn't. Maybe the IFO Overview needs to show the schematic in a way where this doesn't happen: "No light from any laser entering the MC but it still is locked with a resonating field inside."
- Shutting IMCR shutter (hoping that would unlock the IMC), still nothing happend.
- Tried shutting PSL shutter from Rossa, nothing happened to MC lock still.
- Closed shutter IOO>Lock MC> Close PSL and this unlocked the IMC. Found out that this shutter channel is C1:PSL-PSL_ShutterRqst while the one from the sitemap>Shutter>PSL changes C1:AUX-PSL_ShutterRqst. Some clarification on these medm screens would be nice.
- Disabled the MC autolocked from IOO>Lock MC screen (C1:IOO-MC_LOCK_ENABLE).
- Checked the scripts/SUS/freeswing.py to understand how kick is delivered and optic is left to swing freely.
- Next, we are looking at the C1SUS_MC1 screen to understand what channels to read during data acquisition.
- In sensor matrix, we see INMON for each sensor which is probably raw counts data from the OSEMs. Rana mentioned that OSEM data comes out in units of microns. These are C1:SUS-MC1_ULSEN_OUTPUT (and so on for UR, LL, LR, SD).

- In prep for finishing, recovered Autolocker by first opening the PSL mechanical shutter, then re-enabling the Autolocker. The IMC lock didn't immediately recover, and we saw some fuzz on the PSL-FSS_FAST trace, so we closed the shutter again, waited a minute, then re-opened it and MC caught its lock.
 

ELOG V3.1.3-